Tango, Corp. expects to have an earnings per share of $4. The company will be paying out 50% of that earnings to its shareholders, with the rest retained in the company for future growth at rate of 20% each year. The share price of the company's stock is currently at $20.

What rate of return do Tango’s investors require if its stock's intrinsic value has been reflected in the market price? (Do not round intermediate calculations.)

Rate of Return = ?
Answer #1

This question requires application of constant growth dividend discount model according to which: Po - Divi T-9 Po = Price of

Div1 = $4 * 50% = $2

20=- T-02

r - 0.2 = 0.1

r = 0.3

Rate of Return, r = 30.00%